OpenStudy (anonymous):

Simplify [sqrt(36) - 4][sqrt(36) + 4] meaning 6-2 X 6+4=20 is this correct?

OpenStudy (anonymous):

sqrt(36)=6 ----> (6-4)(6+4) = 2*10=20

OpenStudy (anonymous):

(√36 - 4)(√36 + 4) (6 - 4)(6 + 4) (2)(10) 20

OpenStudy (anonymous):

yup but be careful of your solution \[(\sqrt{36}+4)(\sqrt{36}-4)=\sqrt{36}^{2} -4^{2} =36-16=20\]